OpenStudy (anonymous):

f(X)=ln(x)+13x³-11x+cos(x), f'(x)=?

OpenStudy (nubeer):

have u tried solving it?

OpenStudy (anonymous):

no its completly difficult for me

OpenStudy (nubeer):

do you know the derivative of lnx?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

what is it?

OpenStudy (abb0t):

\[f' (\ln(x)) = \frac{ 1 }{ x }\] \[f' (a)^n = na^{n-1}\] You should know the derivative of basic trig functions. They are very common in more advance math and should be well known.

OpenStudy (anonymous):

do u know any link of these basics

OpenStudy (abb0t):

Complete derivatives cheat sheet: http://tutorial.math.lamar.edu/pdf/Calculus_Cheat_Sheet_Derivatives.pdf

OpenStudy (anonymous):

thank u very much